Integrate the following functions:
  1. ∫xsinxcoxdx
  2. ∫ root of tanxdx

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
First is the direct consequence of integration by parts
I = 1/2 \int xsin2x dx = 1/2[-x/2cos2x + 1/4sin2x]
For the second
we put tanx = t^2
then we get on simplifying
I = \int \frac{2t^2}{1+t^4}dt
now this can be done